# Cosmology on the Generalized Proca Theory

### v1

@@ 1.1.2 Energy-Momentum Tensor
$T_{\alpha\beta}=\left[\begin{array}[]{llll}\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}T_{00}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T_{01}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T_{02}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T_{03}\\ \color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T_{10}&\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{11}&\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}T_{12}&\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}T_{13}\\ \color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T_{20}&\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}T_{21}&\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{22}&\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}T_{23}\\ \color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T_{30}&\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}T_{31}&\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}T_{32}&\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{33}\end{array}\right]$ (1.14)
where $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}T_{00}$ is related to the density, $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T_{i0}$ is the energy flux in the $i$ direction, $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T_{0i}$ is the momentum density in the $i$ direction, $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}T_{i(j>i)}$ is the shear stress, $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}T_{i(j<i)}$ is the momentum flux, and $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{ii}$ is related to the pressure. Moreover, if the fluid is isolated, the energy momentum tensor is conserved, i.e., $\nabla_{\mu}T^{\mu\nu}=0$, as is the case for the cosmic fluid, since it represents all the energy content. Additionally, the density and pressure are defined as $\rho=-T^{0}_{0}$ and $p=T^{i}_{i}/3$, respectively, and the relation between these quantities is called the equation of state, $w$, such that
@@
Perfect Fluid,
It is impossible to consider a Universe according to ours without thinking about the energy content present in it. The main assumption is that on large scales, the primary fluids, radiation, matter, and dark matter behave as perfect fluids, i.e., they can be described isoentropically. Thus,
$T^{\mu\nu}=(\rho+p)u^{\mu}u^{\nu}-pg^{\mu\nu},$ (1.16)
where $\rho$ is the fluid density, $p$ the pressure, and $u^{\mu}$ is its four-velocity. Furthermore, the standard fluids can be classified as:
  • Non-relativistic components, such as baryonic matter and cold dark matter (CDM), characterized by not exerting pressure, i.e., their equation of state is $w=0$.
  • Relativistic components, such as electromagnetic radiation, massless neutrinos, and hot dark matter (HDM), characterized by having the equation of state $w=1/3$, as described in the Refs [47, 52].
1.2 The Standard Cosmological Model $\Lambda$CDM
The standard cosmological model, also known as the $\Lambda$CDM model, is founded upon the cosmological principle, which asserts that the Universe is homogeneous and isotropic on large scales. Consequently, the line element is described by Eq. (1.12), and the governing equations at these scales are derived from general relativity, as expressed in Eq. (1.1). Beginning with the description of the fluid in its rest frame, where $u^{i}=0$, it can be demonstrated that the Einstein field equations yield two independent equations, commonly referred to as the Friedman equations:
$3m_{\text{P}}^{2}H^{2}=\rho+m_{\text{P}}^{2}\Lambda-3m_{\text{P}}^{2}\frac{k}{a^{2}}=\rho_{\text{total}}-3m_{\text{P}}^{2}\frac{k}{a^{2}},$ (1.17)
$-2\dot{H}-3H^{2}=\frac{p}{m_{\text{P}}^{2}}+\frac{k}{a^{2}}-\Lambda,$ (1.18)
In Eq. (1.17), the left-hand side represents the geometric aspects, while the right-hand side accounts for the contributions to the density, where $\rho_{\text{total}}=\rho+m_{\text{P}}^{2}\Lambda$, along with the spatial curvature term $3m_{\text{P}}^{2}k^{2}/a^{2}$. The unknowns variables in Eqs. (1.17) and (1.18) are $\rho(t)$, $a(t)$, and $p(t)$; however, $p(t)$ can be determined using the equation of state, Eq. (1.15), and the condition $\nabla_{\mu}T^{\mu\nu}=0$, leading to the fluid continuity equation:
@@ 1.2 The Standard Cosmological Model $\Lambda$CDM
A straightforward method for analyzing such systems is through dynamical analysis. To facilitate this study, it is convenient to introduce dimensionless variables normalized by the critical density $\rho_{c}=3m_{\text{P}}^{2}H^{2}$, resulting in the density parameters for the fluid $\rho$, the cosmological constant, and the curvature parameter:
$\Omega=\frac{\rho}{3m_{\text{P}}^{2}H^{2}},\quad\Omega_{\Lambda}=\frac{\Lambda}{3H^{2}},\quad\Omega_{k}=-\frac{k}{a^{2}H^{2}},$ (1.20)
respectively. The first Friedman equation, Eq. (1.17), then simplifies to
$1=\Omega+\Omega_{\Lambda}+\Omega_{k},$ (1.21)
and the second Friedman equation, Eq. (1.18), can be expressed in terms of the deceleration parameter $q\equiv-\frac{a\ddot{a}}{\dot{a}^{2}}=-1-\frac{\dot{H}}{H^{2}}$:
$q=\frac{1}{2}\Omega(1+3w)-\Omega_{\Lambda}.$ (1.22)
Incorporating $\Omega_{\Lambda}$ into $\Omega$ within the deceleration parameter, we obtain
$q=\frac{1}{2}\Omega(1+3w),$ (1.23)
where it is evident that if $q>0$, the Universe is decelerating, and if $q<0$, the Universe is accelerating. This acceleration occurs when $w<-1/3$, as indicated in Eq. (1.23). Any fluid capable of generating this effect is categorized as dark energy (DE).
1.2.1 Continuity and Energy Decay Equations for Different Types of Fluids in the Universe
The continuity equation for perfect fluids is provided by Eq. (1.19). With knowledge of the equation of state $p(\rho)$, information about the evolution of the fluid can be deduced. The behavior of relativistic, non-relativistic components, and the cosmological constant can be analyzed with respect to the scale factor.
@@ 1.2.1 Continuity and Energy Decay Equations for Different Types of Fluids in the Universe
For relativistic components, the equation of state is given by $w=\frac{1}{3}$. Consequently, the continuity equation becomes
$\dot{\rho}_{r}+4\frac{\dot{a}}{a}\rho_{r}=0,$ (1.24)
with a solution that satisfies
$\rho_{r}\propto\frac{1}{a^{4}},$ (1.25)
indicating that relativistic components decay proportionally to $a^{4}$ as the Universe expands. Additionally, assuming a flat Universe ($k=0$), the differential equation (1.17) can be solved with $\rho_{\text{total}}=\rho_{r}$. Given that $H=\dot{a}/a$, we obtain $a\propto t^{1/2}$.
Non-Relativistic Components
@@
For non-relativistic components, the equation of state is $w=0$. Hence, the continuity equation becomes
$\dot{\rho}_{m}+3\frac{\dot{a}}{a}\rho_{m}=0,$ (1.26)
with a solution that satisfies
$\rho_{m}\propto\frac{1}{a^{3}},$ (1.27)
indicating that non-relativistic components decay as $a^{3}$, corresponding to the volume expansion of the Universe. Furthermore, assuming a flat Universe, the differential equation (1.17) can be solved with $\rho_{\text{total}}=\rho_{m}$, yielding $a\propto t^{2/3}$.
Dark Energy with a Constant Equation of State $w$
For dark energy (DE), characterized by $w<-\frac{1}{3}$, the continuity equation is
$\dot{\rho}_{\text{DE}}+3(1+w)\frac{\dot{a}}{a}\rho_{\text{DE}}=0,$ (1.28)
with a solution that satisfies
$\rho_{\text{DE}}\propto\frac{1}{a^{3(1+w)}},$ (1.29)
indicating that dark energy components decay as $a^{3(1+w)}$ if $w>-1$, remain constant if $w=-1$ (as in the case of $\Lambda$), and grow if $w<-1$ as the Universe expands. Consequently, if this form of dark energy exists, it is expected to eventually dominate the energy content of the Universe, as it increasingly overtakes matter and radiation [^2] Henceforth, the terms ”radiation” and ”relativistic matter” will be used interchangeably, as well ”non-relativistic matter” and ”matter.”. In addition, assuming a flat Universe, the differential equation (1.21) can be solved for $\rho_{\text{total}}=m_{\text{P}}^{2}\Lambda$, i.e., $w=-1\rightarrow a(t)\propto e^{t\sqrt{\Lambda/3}}$, or for $\rho_{\text{total}}=\rho_{\text{DE}}$, yielding $w\neq-1\rightarrow a(t)\propto t^{2/(3(1+w))}$ [52].
1.3 Problems of the Standard Cosmological Model
1.3.1 Flatness Problem
The flatness problem can be seen from the equation Eq. (1.21) by taking $\Omega_{\text{total}}=\Omega+\Omega_{\Lambda}$, such that,
$|\Omega_{\text{total}}(t)-1|=|\Omega_{k}(t)|,$ (1.30)
if $\Omega_{k}(t_{i})=0$ it is because $k=0$ as $a(t_{i})H(t_{i})<\infty$, Eq. (1.20), since these are measurable physical quantities, therefore, $\Omega_{k}(t)=0$ for all $t$, meaning that if the Universe starts flat, it remains flat. Additionally, if we study the behavior of $a^{2}H^{2}$ during the early evolution of the Universe, assuming that only radiation or matter is present, we find that
  • For radiation dominance
    $a^{2}H^{2}\propto t^{-1}\Longrightarrow\left|\Omega_{\text{total}}(t)-1\right|\propto t.$ (1.31)
  • For matter dominance
    $a^{2}H^{2}\propto t^{-2/3}\Longrightarrow\left|\Omega_{\text{total}}(t)-1\right|\propto t^{2/3}.$ (1.32)
[…]
The flatness problem arises from the lack of a theoretical mechanism to generate these initial conditions $\left|\Omega_{\text{total}}(t_{0})-1\right|\approx 0$. In other words, if $\left|\Omega_{\text{total}}(t_{0})-1\right|$ is an increasing function of time, what mechanism causes $\Omega_{\text{total}}(t_{0})$ to be very close to 1 today? The only remaining solution is to make what is known as a fine-tuning of the initial conditions and assume that the Universe was always flat. However, these ideas are not very physically acceptable.
1.3.2 Horizon Problem
[…]
Figure 1.5: Schematic of the observational horizon of the Universe. The center of the semicircle represents our location in the Universe, points $A$ and $B$ are two regions located at the antipodes of our causally disconnected Universe since they are not within their observational horizons.
1.3.3 Open Problems
The attempt to study the current expansion of the Universe and measure the Hubble constant today $H_{0}$ has represented one of the greatest challenges in physics of our time. There are many ways to obtain it. One way is model-dependent, usually the $\Lambda$CDM model based on data detected by the PLANCK project ($H_{0}=(67.27\pm 0.60)k\text{m}/\text{s}Mpc$ with $68\%$ confidence)[1, 12, 53], and the other way is based on measuring local phenomena, such as those done by the SH0ES collaboration ($H_{0}=(73.04\pm 1.04)k\text{m}/\text{s}Mpc$ with $68\%$ confidence)[54, 55, 53]. The main problem here is not that there are various ways to measure $H_{0}$, but that the measurements based on the CMB and local measurements based on supernovae Ia have a discrepancy of around $5\sigma$, known as the $\boldsymbol{H_{0}}$ tension[56, 18]. Besides the $H_{0}$ tension, there is also the cosmological constant problem, as there is no fundamental explanation for introducing the cosmological constant $\Lambda$. Initially, the most natural correspondence was between $\Lambda$ and the quantum vacuum energy. However, when comparing observational results with the predictions of the standard model of particle physics, a difference of 120 orders of magnitude is found[57, 58]. Additionally, observations have also shown other possible cosmological anomalies, indicating that the $\Lambda$CDM model may need modification, considering dynamic dark energy models or corrections in the geometry of the Universe, as it has been shown that such models could slightly correct these problems[53].
1.4 Inflationary Universe
In 1981, Alan Guth proposed inflation as a solution to the horizon and flatness problems[59, 52]. Inflation is defined as a period in the Universe during which its expansion occurs in an accelerated manner,
$\text{INFLATION}\Longleftrightarrow\ddot{a}(t)>0,$ (1.33)
as seen in Eq. (1.23). To produce accelerated expansion, it is necessary that the equation of state of the fluid satisfies the relation $w<-1/3$. The typical example is $\Lambda$, as it behaves like a fluid with equation of state $w=-1$ and causes the scale factor to grow as
$a(t)=e^{\sqrt{\frac{\Lambda}{3}}t},$ (1.34)
that is, when the dominance is the cosmological constant, it causes the expansion of the Universe to occur in an accelerated manner.
1.4.1 Solution to the Flatness Problem
Starting from Eq. (1.30), it was observed that in the dominances of radiation and matter, the spatial curvature of the Universe always increases, Eqs. (1.31) and (1.32). When a dark energy-dominated epoch based on $\Lambda$ is added, it is possible to reduce this effect because
$\ddot{a}>0\quad\Longrightarrow\quad\frac{\mathrm{d}}{\mathrm{d}t}(\dot{a})>0\quad\begin{array}[]{c}\Longrightarrow\\ H=\dot{a}/a\end{array}\quad\frac{\mathrm{d}}{\mathrm{d}t}(aH)>0.$ (1.35)
Therefore, the presence of the cosmological constant drastically brings the conditions closer to flatness,
$|\Omega_{\text{total}}(t)-1|\propto\exp{\left(-\sqrt{\frac{4}{3}\Lambda}t\right)}.$ (1.36)
However, although the inclusion of this fluid in the early stage of the Universe sets the initial conditions very close to $\Omega_{\text{total}}\sim 1$, it will not remain flat throughout its evolution. In the future, it may tend to deviate from a flat Universe, as illustrated in Figure 1.6.

### v2

@@ 1.1.4 The Cosmic Fluid: Energy-Momentum Tensor and Equations of State
$T_{\alpha\beta}=\left[\begin{array}[]{cccc}\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}T_{00}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T_{01}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T_{02}&\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T_{03}\\ \color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T_{10}&\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{11}&\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}T_{12}&\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}T_{13}\\ \color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T_{20}&\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}T_{21}&\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{22}&\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}T_{23}\\ \color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T_{30}&\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}T_{31}&\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}T_{32}&\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{33}\end{array}\right]$ (1.16)
Specifically:
  • $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}T_{00}$ is the energy density.
  • $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T_{0i}=\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T_{i0}$ is the energy flux across the $x^{i}$ surface, equivalent to the momentum density in the $i$-th component.
  • $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{ii}$ represents normal stress in the $i$-th coordinate direction, which is called “pressure” when it is the same in every direction, $x^{i}$.
  • $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}T_{ij}$ ($i\neq j$) represents the shear stress, or the flux of the $i$-th component of momentum across the $x^{j}$ surface.
For a closed system like the Universe as a whole, the energy-momentum tensor is conserved, satisfying the covariant conservation equation $\nabla_{\mu}T^{\mu\nu}=0$. This equation leads to the fundamental laws governing the evolution of energy and momentum density in an expanding universe.
The Perfect Fluid Approximation
In cosmology, the large-scale distribution of matter and radiation is remarkably well-approximated by a perfect fluid. A perfect fluid is characterized by being isotropic in its rest frame, meaning it has no viscosity (shear stress) and no heat conduction. Its energy-momentum tensor takes a simpler form, completely determined by its energy density $\rho$ and isotropic pressure $p$:
$T^{\mu\nu}=(\rho+p)u^{\mu}u^{\nu}+pg^{\mu\nu},$ (1.17)
where $u^{\mu}$ is the fluid’s four-velocity ($u^{\mu}u_{\mu}=-1$), which in the comoving frame of the FLRW metric is simply $u^{\mu}=(1,0,0,0)$, and $g^{\mu\nu}$ is the metric tensor (from Eq. (1.15)). In this frame, $T^{00}=\rho$ and $T^{i}_{j}=p\,\delta^{i}_{j}$.
The thermodynamic properties of the cosmic fluid components are often summarized by their equation of state, which relates pressure to energy density, typically expressed as a constant parameter $w$:
@@
The primary components making up the cosmic inventory within the standard $\Lambda$CDM model are modeled as perfect fluids with distinct equations of state:
  • Non-relativistic Matter: This includes baryonic matter and cold dark matter (CDM). These components have negligible random velocities compared to their rest mass energy, exerting negligible pressure. Thus, they are characterized by $w_{m}=0$.
  • Radiation: This includes photons (like the CMB) and relativistic neutrinos. These particles move at or near the speed of light and exert significant pressure. Relativistic kinematic theory shows $p_{r}=\rho_{r}/3$, so they are characterized by $w_{r}=1/3$. (See e.g., [5, 6]).
  • Dark Energy/Cosmological Constant: The component driving the observed accelerated expansion behaves like a fluid with negative pressure. A cosmological constant $\Lambda$ corresponds to a fluid with $w_{\Lambda}=-1$. More general dark energy models allow for different, possibly time-varying, $w$.
Understanding these components and their equations of state is essential for applying Einstein’s equations to determine the expansion history of the Universe.
1.2 The Standard Cosmological Model: $\Lambda$CDM Dynamics
The standard cosmological model, commonly known as the $\Lambda$CDM model, rests upon the cosmological principle. Consequently, the geometry of spacetime is described by the FLRW line element, Eq. (1.15), and the governing dynamics at cosmological scales are dictated by Einstein’s general relativity, expressed in Eq. (1.1). By considering the cosmic medium as a perfect fluid in its rest frame ($u^{i}=0$), the Einstein field equations yield two fundamental dynamical relations, often referred to as the Friedmann equations:
$3m_{\text{P}}^{2}H^{2}$ $=\rho_{\text{total}}-3m_{\text{P}}^{2}\frac{k}{a^{2}},$ (1.19)
$-2m_{\text{P}}^{2}\dot{H}-3m_{\text{P}}^{2}H^{2}$ $=p+m_{\text{P}}^{2}\frac{k}{a^{2}}-m_{\text{P}}^{2}\Lambda.$ (1.20)
In the first Friedmann equation (1.19), the left-hand side represents the geometric aspect, while the right-hand side incorporates the total energy of the sources (where $\rho_{\text{total}}=\rho_{m}+\rho_{r}+m_{\text{P}}^{2}\Lambda$), and the spatial curvature term proportional to $k/a^{2}$. The primary unknown functions in Eqs. (1.19) and (1.20) are the scale factor $a(t)$ (via $H$ and $\dot{H}$), the density $\rho(t)$, and the pressure $p(t)$. The pressure $p(t)$ is related to the density via the equation of state, $p=w\rho$ (Eq. (1.18)). Furthermore, the requirement of energy-momentum conservation, $\nabla_{\mu}T^{\mu\nu}=0$, provides the fluid continuity equation, which governs the evolution of the density:
@@ 1.2 The Standard Cosmological Model: $\Lambda$CDM Dynamics
For analyzing the cosmic dynamics, it is often convenient to introduce dimensionless density parameters by normalizing the energy densities by the factor $3m_{\text{P}}^{2}H^{2}$. We define the density parameters for the fluid component $\rho$, the cosmological constant $\Lambda$, and the spatial curvature $k$ as:
$\Omega=\frac{\rho}{3m_{\text{P}}^{2}H^{2}},\quad\Omega_{\Lambda}=\frac{\Lambda}{3H^{2}}\quad\text{ and }\quad\Omega_{k}=-\frac{k}{a^{2}H^{2}},$ (1.22)
respectively [^1] Note that $\Omega$ here refers specifically to the matter/radiation fluid density parameter, distinct from $\Omega_{\Lambda}$.. With these definitions, the first Friedman equation (1.19) simplifies to the well-known constraint equation:
$1=\Omega+\Omega_{\Lambda}+\Omega_{k}.$ (1.23)
The second Friedman equation (1.20) can be related to the cosmic acceleration via the deceleration parameter $q\equiv-a\ddot{a}/\dot{a}^{2}=-(1+\dot{H}/H^{2})$. Expressed in terms of the density parameters, it yields:
$q=\frac{1}{2}\sum_{i}\Omega_{i}(1+3w_{i})=\frac{1}{2}\Omega_{\text{eff}}(1+3w_{\text{eff}}),$ (1.24)
where the effective density parameter is $\Omega_{\text{eff}}=\Omega+\Omega_{\Lambda}$ and the effective equation of state $w_{\text{eff}}$ satisfies $p_{\text{total}}=w_{\text{eff}}\rho_{\text{total}}$. From either perspective, accelerated expansion ($q<0$) requires a significant contribution from a component with negative pressure. For a single dominant fluid with equation of state $w$, acceleration occurs when $w<-1/3$. Any fluid capable of generating this effect ($q<0$) is broadly categorized as dark energy (DE).
1.2.1 Continuity and Energy Decay Equations for Different Cosmic Fluids
The continuity equation (1.21) dictates the evolution of the energy density $\rho_{i}$ for any component with a specific equation of state $w_{i}=p_{i}/\rho_{i}$. Let us analyze the behavior of the key components within $\Lambda$CDM as a function of the scale factor $a$, assuming a spatially flat ($k=0$) universe for simplicity in deriving the $a(t)$ dependencies.
@@ 1.2.1 Continuity and Energy Decay Equations for Different Cosmic Fluids
Relativistic components, such as photons and neutrinos, are characterized by $w_{r}=1/3$. The continuity equation (1.21) becomes:
$\dot{\rho}_{r}+3H(1+1/3)\rho_{r}=\dot{\rho}_{r}+4H\rho_{r}=0.$ (1.25)
Integrating this equation yields the solution:
$\rho_{r}\propto a^{-4}.$ (1.26)
Thus, the energy density of relativistic components decays rapidly as $\propto a^{-4}$ during cosmic expansion. If radiation dominates the energy budget ($\rho_{\text{total}}\approx\rho_{r}$), the first Friedman equation (1.19) implies $H^{2}\propto a^{-4}$, leading to $a(t)\propto t^{1/2}$.
Non-Relativistic Components (Matter)
Non-relativistic components (baryons and cold dark matter) exert negligible pressure, so $w_{m}=0$. The continuity equation then reads:
$\dot{\rho}_{m}+3H(1+0)\rho_{m}=\dot{\rho}_{m}+3H\rho_{m}=0.$ (1.27)
The solution is:
$\rho_{m}\propto a^{-3}.$ (1.28)
This indicates that the energy density of non-relativistic matter dilutes proportionally to the volume ($a^{3}$) as the Universe expands. If matter dominates, Eq. (1.19) implies $H^{2}\propto a^{-3}$, resulting in $a(t)\propto t^{2/3}$.
Dark Energy with a Constant Equation of State $w$ ($w$CDM)
Dark energy broadly refers to any component capable of driving cosmic acceleration ($q<0$), requiring $w<-1/3$. The simplest model beyond the cosmological constant ($\Lambda$, where $w=-1$) is one where dark energy has a constant equation of state parameter $w$. For such a component, the continuity equation is:
$\dot{\rho}_{\text{DE}}+3H(1+w)\rho_{\text{DE}}=0.$ (1.29)
The solution is:
$\rho_{\text{DE}}\propto a^{-3(1+w)}.$ (1.30)
This shows that the energy density evolution depends critically on $w$:
  • If $w>-1$, $\rho_{\text{DE}}$ decreases as the Universe expands (though slower than matter if $w<0$).
  • If $w=-1$ (the $\Lambda$CDM case), $\rho_{\text{DE}}=\rho_{\Lambda}$ remains constant.
  • If $w<-1$ (phantom energy), $\rho_{\text{DE}}$ increases as the Universe expands.
Consequently, if a dark energy component with $w<0$ exists, it dilutes slower than matter and radiation and is expected to eventually dominate the energy content of the Universe [^2] Henceforth, the terms “radiation” and “relativistic matter” will be used interchangeably, as well “non-relativistic matter” and “matter”.. Furthermore, solving the first Friedman equation (1.19) under the dominance of such a component yields specific expansion laws: for $w=-1$, $a(t)\propto e^{Ht}$ where $H=\sqrt{\Lambda/3}$; for $w\neq-1$, $a(t)\propto t^{2/(3(1+w))}$ [6].
1.3 Challenges for the Standard Cosmological Model
Despite the remarkable success of the $\Lambda$CDM model in describing a wide array of cosmological observations, it faces several significant challenges. These include conceptual fine-tuning issues related to the initial conditions of the Universe, as well as emerging tensions between different observational datasets [11, 12, 7, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Such discrepancies hint that $\Lambda$CDM, while a powerful effective theory, might not represent the complete picture of cosmic evolution, potentially requiring extensions involving new physics. Alternative frameworks, such as modified gravity theories or models incorporating dynamical DE, aim to address these outstanding issues  [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43]. In the following subsections, we outline the key problems that motivate the exploration of these alternative approaches.
1.3.1 The Flatness Problem
The flatness problem originates from the relationship between the total energy density and the spatial curvature of the Universe, as expressed by the first Friedmann equation in terms of density parameters, Eq. (1.23). Defining the total density parameter excluding curvature as $\Omega_{\text{total}}=\Omega+\Omega_{\Lambda}$ (where $\Omega$ represents matter and radiation), the equation becomes:
$|\Omega_{\text{total}}(t)-1|=|\Omega_{k}(t)|=\left|-\frac{k}{a^{2}(t)H^{2}(t)}\right|.$ (1.31)
If the Universe begins perfectly flat ($k=0$), then $\Omega_{k}(t_{i})=0$, implying $\Omega_{\text{total}}(t_{i})=1$. Since $a(t)H(t)$ remains finite, this condition $k=0$ ensures $\Omega_{k}(t)=0$ for all times; a flat universe remains flat. However, if there is any initial deviation from flatness ($k\neq 0$), the quantity $a^{2}H^{2}$ evolves differently depending on the dominant energy component. During the radiation and matter-dominated eras, specifically:
  • For radiation dominance ($a\propto t^{1/2}$, $H\propto t^{-1}$):
    $a^{2}H^{2}\propto t^{-1}\quad\Longrightarrow\quad|\Omega_{\text{total}}(t)-1|=|\Omega_{k}(t)|\propto\frac{1}{a^{2}H^{2}}\propto t.$ (1.32)
  • For matter dominance ($a\propto t^{2/3}$, $H\propto t^{-1}$):
    $a^{2}H^{2}\propto t^{-2/3}\quad\Longrightarrow\quad|\Omega_{\text{total}}(t)-1|=|\Omega_{k}(t)|\propto\frac{1}{a^{2}H^{2}}\propto t^{2/3}.$ (1.33)
In both scenarios, the deviation from flatness, $|\Omega_{\text{total}}(t)-1|$, grows with time. This implies that $\Omega_{\text{total}}=1$ is an unstable equilibrium. For the Universe to be observed as nearly flat today ($\Omega_{\text{total}}(t_{0})\approx 1$, implying $|\Omega_{k}(t_{0})|\ll 1$ [7]), its initial value must have been extraordinarily close to unity. For example, during Big Bang Nucleosynthesis (BBN) [^3] Nucleosynthesis is the early epoch in the Universe, roughly $t\sim 1$ s to minutes, where nuclear reactions formed the first light elements as photons had energies around $\sim 1$ MeV., the total density parameter must have satisfied $|\Omega_{\text{total}}(t_{\text{BBN}})-1|\lesssim 10^{-16}$ [6].
The essence of the flatness problem is the lack of a compelling mechanism within the standard Big Bang framework (without inflation) to explain why the initial conditions were so precisely fine-tuned to produce the near-flatness observed today. Why did the Universe start with $\Omega_{\text{total}}$ so incredibly close to 1?
1.3.2 The Horizon Problem
The horizon problem arises from the remarkable large-scale isotropy observed in the Cosmic Microwave Background (CMB). The CMB radiation, emitted roughly 380,000 years after the Big Bang, exhibits an almost perfectly uniform temperature of $T_{0}\approx 2.725\,\text{K}$ across the entire sky, with fluctuations only at the level of $\Delta T/T\sim 10^{-5}$ [7].
In the standard Big Bang model (without inflation), the Universe has a finite age, implying a finite particle horizon – the maximum distance light could have traveled since the beginning. Calculations show that at the time of CMB emission (last scattering), regions on the sky separated by more than $\sim 1^{\circ}$ were outside each other’s particle horizons[5]. They were causally disconnected, meaning they could not have exchanged signals or interacted to reach thermal equilibrium. Yet, these causally separated regions are observed to have the same temperature to extremely high precision. How did regions that never communicated achieve such thermal homogeneity? This conundrum is illustrated schematically in Figure 1.4.
Figure 1.4: Schematic illustration of the horizon problem. An observer (center) sees CMB radiation from points A and B, which are widely separated on the surface of last scattering. In the standard Big Bang model without inflation, the particle horizons of A and B (dashed circles) do not overlap, meaning they were causally disconnected when the CMB was emitted, yet they exhibit nearly identical temperatures.
1.3.3 Inflationary Solution to Primordial Puzzles
The flatness and horizon problems find an elegant resolution within the paradigm of cosmic inflation, proposed by Guth and Linde [44, 6]. Inflation postulates a period of quasi-exponential, accelerated expansion ($\ddot{a}>0$) in the very early Universe. As established earlier (see Eq. (1.24)), such acceleration requires a dominant component with an equation of state $w<-1/3$. A simple example is cosmological constant $\Lambda$, which drives $a(t)\propto e^{\sqrt{\Lambda/3}t}$.
Inflation addresses the primordial puzzles as follows:
  • Solving the Flatness Problem: During a period of accelerated expansion, $\ddot{a}>0$. This implies $\frac{d}{dt}(aH)=\frac{d}{dt}(\dot{a})>0$. Consequently, the denominator $a^{2}H^{2}$ in the expression for $|\Omega_{k}(t)|=|k|/(a^{2}H^{2})$ grows enormously. Any initial curvature is stretched to near-zero values, driving $|\Omega_{\text{total}}(t)-1|$ exponentially towards zero. Inflation dynamically forces the Universe towards flatness, alleviating the need for extreme fine-tuning of initial conditions (see Figure 1.5).
  • Solving the Horizon Problem: Prior to inflation, the region corresponding to our entire observable Universe today was microscopically small and causally connected, allowing it to reach thermal equilibrium. Inflation then stretched this tiny, homogeneous patch by an enormous factor (at least $e^{60}\approx 10^{26}$), making it encompass scales vastly larger than the post-inflationary particle horizon. Thus, the large-scale uniformity of the CMB reflects the equilibrium established before inflation expanded the region beyond causal contact (see Figure 1.6).
  • Seeding Structure: Beyond solving these problems, inflation also provides a natural mechanism for generating the primordial density fluctuations needed to seed galaxy formation. Quantum fluctuations inherent in the inflating field are stretched to cosmological scales, creating the nearly scale-invariant, adiabatic, Gaussian perturbations observed in the CMB and large-scale structure [8].